OpenStudy (anonymous):

SERIOUSLY NEED HELP!! PLEASE!! PQRS is an inscribed quadrilateral whose diagonals intersect at T. Segment PQ is parallel to segment SR, as shown below. *Prove that if angle SQP is 55° and angle QSP is 40°, then angle SPR is 30°. Write a two-column proof showing statements and reasons.

OpenStudy (anonymous):

Does this help

jimthompson5910 (jim_thompson5910):

one moment, I'm trying to see how to connect finding angle R with finding angle SPR

jimthompson5910 (jim_thompson5910):

angle QSR is congruent to angle SQP since they are alternate interior angles So angle QSR = 55 ------------------ Angles RQP and RSP add to 180 (since they are part of an inscribed quadrilateral) So RPQ + RSP = 180 RPQ + (55+40) = 180 RPQ + 95 = 180 RPQ = 180-95 RPQ = 85 So angle SQR is 30 degrees (since 85 - 55 = 30)

OpenStudy (anonymous):

Ok thanks